Solve the following equations. 
 (7y + 4)/(y + 2) = -4/3 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

(7y+4)/(y+2)=-4/3
⇒7y+4=-4/3(y+2)
⇒3(7y+4)=-4(y+2)
⇒21y+12=-4y-8
⇒21y+4y=-8-12
⇒25y=-20
⇒y=-20/25=-4/5

Thanks